# Tubes in sub-Riemannian geometry and a Weyl’s invariance result for curves in the Heisenberg groups

### v2

@@ 1. Introduction
Let $S$ be a closed submanifold of codimension $m$ in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$. In [Wey39], Weyl derived a formula for the volume of a tube of small radius $r$ around $S$:
@@
Inspired by Weyl’s results, we study the volume of tubes in more general metric spaces, and more precisely in (sub-)Riemannian manifolds. This research is motivated by the recent works in this area [Fer07, BFF+15, Rit21], which focused on the case of tubes around hypersurfaces embedded in the Heisenberg group, and [BB24], for tubes around hypersurfaces in three-dimensional contact structures. (In this case, the corresponding volume asymtptotics are known as Steiner’s formulas.) See also [BBL20] where, with a different spirit, the authors study the volume of small balls in three-dimensional contact structures.
We develop the theory at the greatest level of generality (in the $C^{k}$ category), addressing the case of non-characteristic submanifolds of arbitrary codimension embedded in a general, possibly rank-varying, sub-Riemannian manifold. In the specific case of the Heisenberg groups, we prove that the volume of small tubes around curves does not depend on the way the curve is isometrically embedded, but only on its Reeb angle. This result extends the classical Euclidean one due to Weyl. Its proof is based on a symmetry argument of independent interest, new even in the Euclidean setting.
@@ 1.1. Regularity of the distance and tubular neighbourhoods
The study of tubes begins with the definition of what a tube is, and instrumental to this is the study of the regularity of the distance from a submanifold. This constitutes our first result, which continues the research line initiated in [AF07, AF08, Fer07, Rit21] for $C^{2}$ hypersurfaces or submanifolds of the Heisenberg groups and [AFM17] for $C^{\infty}$ hypersurfaces of some special step $2$ Carnot groups. Our contribution extends all previous results to $C^{k}$ submanifolds of arbitrary codimension, embedded in general rank-varying sub-Riemannian manifolds, with optimal regularity in the $C^{k}$ class.
We refer to Section 2 for all relevant definitions. In the following, $\delta:M\to\mathbb{R}$ is the distance from an embedded submanifold $S\subset M$, $AS\subset T^{*}M$ is the bundle of covectors that vanish on $TS$ (the so-called annihilator bundle), $E:AS\to M$ is the restriction of the sub-Riemannian exponential map to $AS$, $H\in C^{\infty}(T^{*}M)$ denotes the sub-Riemannian Hamiltonian, while $\nabla$ represents the horizontal gradient of a $C^{1}$ function. See Theorem 3.1.
@@
Without adding further hypotheses, the assumption $k\geq 2$ is necessary, already in the Euclidean case, see [KP81, Foo84]. The non-characteristic assumption is also crucial: $\delta$ is not even Lipschitz in charts at characteristic points, see [ACS18].
@@
Items i, LABEL: and ii immediately imply Items iii, LABEL: and iv with $C^{k-1}$ regularity. The improvement to $C^{k}$ regularity, $k\geq 2$, requires more work. This unexpected gain of regularity was observed in the Euclidean case in [KP81, Foo84].
@@
A version of Theorem 1.1 appeared in [Ros24] when $S$ is closed and smooth ($k=\infty$), building on the codimension $1$ case treated in [FPR20]. With respect to those references, Items iii, LABEL:, iv, LABEL: and v are new and require new arguments for $2\leq k<+\infty$. In [ACS18], among other results, a local analogue of Items iii and iv is obtained for smooth hypersurfaces ($k=\infty$, $m=1$), with different techniques. A version of Theorem 1.1 for the Heisenberg groups was proven in [Rit21], using the explicit knowledge of minimizing geodesics.
@@ 1.2. The volume of tubes
Our second main result is on the asymptotics of the volume of tubes. We were inspired, in particular, by the approach of [BFF+15], who introduced the concept of iterated divergences, to account for the lack of a tensorial calculus well-adapted to the sub-Riemannian setting, akin to the one used in deriving the classical Weyl’s tube formula [Gra04]. Here, we improve and extend this approach to general codimension.
@@
Let $M$ be a smooth (or real-analytic) sub-Riemannian manifold, equipped with a smooth (or real-analytic) measure $\mu$. Let $S\subset M$ be a bounded non-characteristic embedded submanifold of codimension $m\geq 1$, of class $C^{2}$ (without boundary) and with the extendibility property. Let $r_{0}=r_{0}(S)>0$ be its injectivity radius. Then the volume of the tube $r\mapsto\mu(T(r))$ is smooth (or real-analytic) on $[0,r_{0})$. Furthermore, there exists a continuous density $\sigma_{m}$ on $A^{1}S$, defined by
@@
In this generality, one cannot expect $\mu(T(r))$ to be a polynomial in $r$, as it happens for Euclidean tubes. However by the above theorem, as soon as the ambient structure is smooth (resp. real-analytic), the volume of tubes is smooth (resp. real-analytic) even if the corresponding submanifold $S$ is just $C^{2}$. Note that the real-analyticity of the volume of the tube around smooth hypersurfaces ($k=\infty$, $m=1$) in the three-dimensional Heisenberg group was obtained in [BFF+15] as a consequence of the explicit computation of all coefficients in the Taylor expansion. Our proof uses different ideas.
If $S$ is a two-sided hypersurface, and one considers its (positive) half-tube $T^{+}(r)$, the density $\sigma_{1}$ and the functions $w_{1}^{(j)}$ of Theorem 1.10 can be identified, respectively, with the sub-Riemannian perimeter measure $\sigma$ and the iterated divergences ${\rm div}^{j}_{\mu}(\nabla\delta_{\mathrm{sign}})$ on $S$. In this setting, we generalize the Steiner’s formula obtained in [BFF+15]. See Theorem 4.2.
@@
In the setting of Theorem 1.10, assume furthermore that $S$ is a two-sided hypersurface. Then the volume of the half-tube $r\mapsto\mu(T^{+}(r))$ is smooth (or real-analytic) on $[0,r_{0})$. Moreover, $\mu(T^{+}(r))$ has the following Taylor expansion at $r=0$:
@@ 1.3. Polynomial character of iterated divergences
where $F_{1},\dots,F_{4}$ are defined as above replacing $\delta$ with $\delta_{\mathrm{sign}}$. Compare with [BFF+15], where the authors expressed the integrands in terms of $5$ basic functional invariants: $F_{1},\dots,F_{4}$ and, in addition, $F_{5}:=(X_{0}X_{1}\delta)^{2}+(X_{0}X_{2}\delta)^{2}$. However, by using the Eikonal equation and its derivative, one can show that $F_{5}=F_{2}^{2}$. We stress that, in order to prove 1.14, we do not compute explicitly the iterated divergences.
@@ 1.4. Weyl’s invariance for tubes around curves of the Heisenberg groups
One of the deep results in [Wey39] is that the coefficients appearing in the tube formula have intrinsic nature. This can be stated as follows: any isometric embedding of a given Riemannian manifold $S$ in $\mathbb{R}^{n}$ has the same tube volume for small enough radii. Broadly speaking, we call this result Weyl’s invariance. To our best knowledge, all proofs of this fact pass through the computation of the Jacobian determinant of the normal exponential map, explicit spherical integration, and Gauss’ formula (relating the second fundamental form of an Euclidean submanifold with its Riemannian curvature), in order to cast the coefficients in terms of the Riemannian curvature of $S$. See [Gra04]. We remark the following fact about Weyl’s invariance result:
@@
We can now state the invariance result. The corresponding statement for curves in the Euclidean space is due to Hotelling [Hot39], which motivated [Wey39]. We remark that this is the first time that such a result is obtained in sub-Riemannian geometry. (cf. Theorem 6.3).
@@ 2. Preliminaries
We recall some basic notions of sub-Riemannian geometry, cf. [ABB20].
@@ 2.1. Sub-Riemannian structures
Let $M$ be a smooth, connected $n$-dimensional manifold. A smooth sub-Riemannian structure on $M$ is defined by generating family of $N\in\mathbb{N}$ global smooth vector fields $\{X_{1},\ldots,X_{N}\}$. The latter defines a (possibly rank-varying) distribution $\mathcal{D}$:
@@ 2.2. Horizontal gradient
(See [RS23, Appendix A] for a proof in the rank-varying case.)
@@ 2.3. Geodesics and Hamiltonian flow
where $\{X_{1},\ldots,X_{N}\}$ is a generating family for the sub-Riemannian structure, and $\langle\lambda,\cdot\rangle$ denotes the action of the covector $\lambda$ on vectors. The Hamiltonian vector field $\vec{H}$ on $T^{*}M$ is then defined by $\varsigma(\cdot,\vec{H})=dH$, where $\varsigma\in\Lambda^{2}(T^{*}M)$ is the canonical symplectic form.
@@ 2.4. Geodesics from a submanifold
If $\gamma:[0,1]\to M$ is a minimizing geodesic from $S$ there exists a corresponding extremal lift, $\lambda:[0,1]\to T^{*}M$, must satisfy the transversality conditions of [AS04, Thm. 12.4], namely
@@ 3. Regularity of the distance and tubular neighbourhoods
The proof of Theorem 3.1 is split in two parts. In the first part, we prove Items i and ii. This part is quite standard, and it extends the proofs contained in [Ros24, FPR20] to the case of a non-smooth, non-compact $S$. We present it here in order mainly to provide a self-consistent argument. From this, Items iii and iv with $C^{k-1}$ regularity in place of $C^{k}$ follow. This is the starting point for the second part of the proof, where some new ideas are needed. There, we fully prove $C^{k}$ regularity in Items iii and iv, and we prove Item v.
@@
The fact that the Eikonal equation (45) holds is classical. Indeed, since $\delta$ is $1$-Lipschitz w.r.t. ${\sf d}$, then (see e.g. [FHK99, Thm. 8]), it holds
@@
We also recall [Rif14, Lemma 2.15]. There, the sub-Riemannian structure is assumed to have constant rank, but the statement holds, with the same proof, in our setting.
@@ 3.2. Uniform tubular neighbourhoods
Any closed manifold $S$ (i.e., compact without boundary) embedded in $M$ satisfies the extendibility property with $\tilde{S}=S$. If $\mathrm{cl}(S)\subset M$ is an embedded submanifold with boundary, then one can build $\tilde{S}$ such that (76) is satisfied, via a suitable flowout from its boundary; if $\mathrm{cl}(S)$ has no characteristic points, $\tilde{S}$ can be chosen to satisfy the same property. The localizing neighbourhoods used in [BFF+15, BB24] also satisfy 3.6. Note that $S$ may satisfy the extendibility property without $\mathrm{cl}(S)$ being an embedded manifold with boundary (e.g., any open set $S\subset\mathbb{R}^{2}\hookrightarrow\mathbb{R}^{3}$, with irregular frontier). An example without the extendibility property is the figure $\infty$ with the central point removed, embedded in $\mathbb{R}^{2}$.
@@ 4. The volume of tubes
Let $M$ be a smooth (or real-analytic) sub-Riemannian manifold, equipped with a smooth (or real-analytic) measure $\mu$. Let $S\subset M$ be a bounded non-characteristic embedded submanifold of codimension $m\geq 1$, of class $C^{2}$ (without boundary) and with the extendibility property. Let $r_{0}=r_{0}(S)>0$ be its injectivity radius. Then the volume of the tube $r\mapsto\mu(T(r))$ is smooth (or real-analytic) on $[0,r_{0})$. Furthermore, there exists a continuous density $\sigma_{m}$ on $A^{1}S$, defined by
@@
Let $\tilde{S}$ be the extension coming from the extendibility property, see 3.6. Thanks to the latter, we can cover $S$ with a finite number of charts where $S$ is the graph of a $C^{2}$ function. Thus, by a partition of unit argument, it is sufficient to consider the case in which $\mathrm{cl}(S)$ is contained in a single chart and is diffeomorphic to the closure of an open set of $\mathbb{R}^{n-m}$. More precisely, we assume there are smooth (or real-analytic) coordinates $(x,y)\in\mathbb{R}^{n-m}\times\mathbb{R}^{m}$ on an open neighbourhood $\mathcal{O}\subset M$ such that $\mathrm{cl}(S)\subset\mathcal{O}$, and there exists a $C^{2}$ function $h:\mathbb{R}^{n-m}\to\mathbb{R}^{m}$ and an open set $B\subset\mathbb{R}^{n-m}$ such that
@@
From the claim, it follows that $r\mapsto\mu(T(r))$ is smooth (or real-analytic) on $[0,r_{0})$, and one can differentiate under the integral sign an infinite number of times in (86).
@@
so that the right hand side of (92) vanishes if $k-m$ is odd, upon integration on $\mathbb{S}^{m-1}$.
@@
In the setting of Theorem 4.1, assume furthermore that $S$ is a two-sided hypersurface. Then the volume of the half-tube $r\mapsto\mu(T^{+}(r))$ is smooth (or real-analytic) on $[0,r_{0})$. Moreover, $\mu(T^{+}(r))$ has the following Taylor expansion at $r=0$:

### v3

@@ 1. Introduction
Let $S$ be a closed submanifold of codimension $m$ in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$. In [26], Weyl derived a formula for the volume of a tube of small radius $r$ around $S$:
@@
Inspired by Weyl’s results, we study the volume of tubes in more general metric spaces, and more precisely in (sub-)Riemannian manifolds. This research is motivated by the recent works in this area [10, 7, 22], which focused on the case of tubes around hypersurfaces embedded in the Heisenberg group, and [9], for tubes around hypersurfaces in three-dimensional contact structures. (In this case, the corresponding volume asymtptotics are known as Steiner’s formulas.) See also [8] where, with a different spirit, the authors study the volume of small balls in three-dimensional contact structures.
We develop the theory for $C^{k}$ non-characteristic submanifolds of arbitrary codimension embedded in a general, possibly rank-varying, sub-Riemannian manifold. In the specific case of the Heisenberg groups, we prove that the volume of small tubes around curves does not depend on the way the curve is isometrically embedded, but only on its Reeb angle. This result extends the classical Euclidean one due to Weyl. Its proof is based on a symmetry argument of independent interest, new even in the Euclidean setting.
@@ 1.1. Regularity of the distance and tubular neighbourhoods
The study of tubes begins with the definition of what a tube is, and instrumental to this is the study of the regularity of the distance from a submanifold. This constitutes our first result, which continues the research line initiated in [4, 5, 10, 22] for $C^{2}$ hypersurfaces or submanifolds of the Heisenberg groups and [6] for $C^{\infty}$ hypersurfaces of some special step $2$ Carnot groups. Our contribution extends all previous results to $C^{k}$ submanifolds of arbitrary codimension, embedded in general rank-varying sub-Riemannian manifolds, with optimal regularity in the $C^{k}$ class.
In the following, $\delta:M\to\mathbb{R}$ is the distance from an embedded submanifold $S\subset M$, $\pi:T^{*}M\to M$ is the cotangent bundle projection, $AS\subset T^{*}M$ is the bundle of covectors that vanish on $TS$ (the so-called annihilator bundle), $E:AS\to M$ is the restriction of the sub-Riemannian exponential map to $AS$, $H\in C^{\infty}(T^{*}M)$ denotes the sub-Riemannian Hamiltonian, while $\nabla$ represents the horizontal gradient of a $C^{1}$ function. We refer to Section 2 for precise definitions. The next result corresponds to Theorem 3.1.
@@
Without adding further hypotheses, the assumption $k\geq 2$ is necessary, already in the Euclidean case, see [17, 11]. The non-characteristic assumption is also crucial: $\delta$ is not even Lipschitz in charts at characteristic points, see [3].
@@
Items i, LABEL: and ii immediately imply Items iii, LABEL: and iv with $C^{k-1}$ regularity. The improvement to $C^{k}$ regularity, $k\geq 2$, requires more work. This unexpected gain of regularity was observed in the Euclidean case in [17, 11].
@@
A version of Theorem 1.1 appeared in [25] when $S$ is closed and smooth ($k=\infty$), building on the codimension $1$ case treated in [12]. With respect to those references, Items iii, LABEL:, iv, LABEL: and v are new and require new arguments for $2\leq k<+\infty$. In [3], among other results, a local analogue of Items iii and iv is obtained for smooth hypersurfaces ($k=\infty$, $m=1$), with different techniques. A version of Theorem 1.1 for the Heisenberg groups was proven in [22], using the explicit knowledge of minimizing geodesics.
@@ 1.2. The volume of tubes
Our second main result is on the asymptotics of the volume of tubes. We were inspired, in particular, by the approach of [7], who introduced the concept of iterated divergences, to account for the lack of a tensorial calculus well-adapted to the sub-Riemannian setting, akin to the one used in deriving the classical Weyl’s tube formula [14]. Here, we improve and extend this approach to general codimension.
@@
Let $M$ be a smooth (or real-analytic) sub-Riemannian manifold, equipped with a smooth (or real-analytic) measure $\mu$. Let $S\subset M$ be a bounded non-characteristic embedded submanifold of codimension $m\geq 1$, of class $C^{2}$ (without boundary) and with the extendibility property. Let $r_{0}=r_{0}(S)>0$ be its injectivity radius. Then the volume of the tube $r\mapsto\mu(T(r))$ is smooth (or real-analytic) on $0,r_{0})$. Furthermore, there exists a continuous density $\sigma_{m}$ on $A^{1}S$, defined by
@@
In this generality, one cannot expect $\mu(T(r))$ to be a polynomial in $r$, as it happens for Euclidean tubes. However by the above theorem, as soon as the ambient structure is smooth (resp. real-analytic), the volume of tubes is smooth (resp. real-analytic) even if the corresponding submanifold $S$ is just $C^{2}$. Note that the real-analyticity of the volume of the tube around smooth hypersurfaces ($k=\infty$, $m=1$) in the three-dimensional Heisenberg group was obtained in [7] as a consequence of the explicit computation of all coefficients in the Taylor expansion. Our proof uses different ideas.
If $S$ is a two-sided hypersurface, and one considers its (positive) half-tube $T^{+}(r)$, the density $\sigma_{1}$ and the functions $w_{1}^{(j)}$ of Theorem 1.10 can be identified, respectively, with the sub-Riemannian perimeter measure $\sigma$ and the iterated divergences ${\rm div}^{j}_{\mu}(\nabla\delta_{\mathrm{sign}})$ on $S$. In this setting, we generalize the Steiner’s formula obtained in [7]. See Theorem 4.2.
@@
In the setting of Theorem 1.10, assume furthermore that $S$ is a two-sided hypersurface. Then the volume of the half-tube $r\mapsto\mu(T^{+}(r))$ is smooth (or real-analytic) on $0,r_{0})$. Moreover, $\mu(T^{+}(r))$ has the following Taylor expansion at $r=0$:
@@ 1.3. Polynomial character of iterated divergences
where $F_{1},\dots,F_{4}$ are defined as above replacing $\delta$ with $\delta_{\mathrm{sign}}$. Compare with [7], where the authors expressed the integrands in terms of $5$ basic functional invariants: $F_{1},\dots,F_{4}$ and, in addition, $F_{5}:=(X_{0}X_{1}\delta)^{2}+(X_{0}X_{2}\delta)^{2}$. However, by using the Eikonal equation and its derivative, one can show that $F_{5}=F_{2}^{2}$. We stress that, in order to prove 1.14, we do not compute explicitly the iterated divergences.
@@ 1.4. Weyl’s invariance for tubes around curves of the Heisenberg groups
One of the deep results in [26] is that the coefficients appearing in the tube formula have intrinsic nature. This can be stated as follows: any isometric embedding of a given Riemannian manifold $S$ in $\mathbb{R}^{n}$ has the same tube volume for small enough radii. Broadly speaking, we call this result Weyl’s invariance. To our best knowledge, all proofs of this fact pass through the computation of the Jacobian determinant of the normal exponential map, explicit spherical integration, and Gauss’ formula (relating the second fundamental form of an Euclidean submanifold with its Riemannian curvature), in order to cast the coefficients in terms of the Riemannian curvature of $S$. See [14]. We remark the following fact about Weyl’s invariance result:
@@
We can now state the invariance result. The corresponding statement for curves in the Euclidean space is due to Hotelling [16], which motivated [26]. We remark that this is the first time that such a result is obtained in sub-Riemannian geometry. (cf. Theorem 6.3).
@@ 2. Preliminaries
We recall some basic notions of sub-Riemannian geometry, cf. [1].
@@ 2.1. Sub-Riemannian structures
Let $M$ be a smooth, connected $n$-dimensional manifold. A smooth sub-Riemannian structure on $M$ is defined by a generating family of $N\in\mathbb{N}$ global smooth vector fields $\{X_{1},\ldots,X_{N}\}$. The latter defines a (possibly rank-varying) distribution $\mathcal{D}$:
@@ 2.2. Horizontal gradient
(See [24, Appendix A] for a proof in the rank-varying case.)
@@ 2.3. Geodesics and Hamiltonian flow
where $\{X_{1},\ldots,X_{N}\}$ is a generating family for the sub-Riemannian structure, and $\langle\lambda,\cdot\rangle$ denotes the action of the covector $\lambda$ on vectors. The Hamiltonian vector field $\vec{H}$ on $T^{*}M$ is then defined by $\varsigma(\cdot,\vec{H})=dH$, where $\varsigma\in\Lambda^{2}(T^{*}M)$ is the canonical symplectic form ($\varsigma=d\tau$, where $\tau\in\Lambda^{1}(T^{*}M)$ is the tautological form).
@@ 2.4. Geodesics from a submanifold
If $\gamma:[0,1]\to M$ is a minimizing geodesic from $S$ there exists a corresponding extremal lift, $\lambda:[0,1]\to T^{*}M$, must satisfy the transversality conditions of [2, Thm. 12.4], namely
@@ 3. Regularity of the distance and tubular neighbourhoods
The proof of Theorem 3.1 is split into two parts. In the first part, we prove Items i and ii. This part is quite standard, and it extends the proofs contained in [25, 12] to the case of a non-smooth, non-compact $S$. We present it here in order mainly to provide a self-consistent argument. From this, Items iii and iv with $C^{k-1}$ regularity in place of $C^{k}$ follow. This is the starting point for the second part of the proof, where some new ideas are needed. There, we fully prove $C^{k}$ regularity in Items iii and iv, and we prove Item v.
@@
The fact that the Eikonal equation (46) holds is classical. Indeed, since $\delta$ is $1$-Lipschitz w.r.t. $\sf d$}, then (see e.g. [13, Thm. 8]), it holds
@@
We also recall [21, Lemma 2.15]. There, the sub-Riemannian structure is assumed to have constant rank, but the statement holds, with the same proof, in our setting.
@@ 3.2. Uniform tubular neighbourhoods
Any closed manifold $S$ (i.e., compact without boundary) embedded in $M$ satisfies the extendibility property with $\tilde{S}=S$. If $\mathrm{cl}(S)\subset M$ is an embedded submanifold with boundary, then one can build $\tilde{S}$ such that (77) is satisfied, via a suitable flowout from its boundary [19, p. 217]; if $\mathrm{cl}(S)$ has no characteristic points, $\tilde{S}$ can be chosen to satisfy the same property. The localizing neighbourhoods used in [7, 9] also satisfy 3.6. Note that $S$ may satisfy the extendibility property without $\mathrm{cl}(S)$ being an embedded manifold with boundary (e.g., any open set $S\subset\mathbb{R}^{2}\hookrightarrow\mathbb{R}^{3}$, with irregular frontier). An example without the extendibility property is the figure $\infty$ with the central point removed, embedded in $\mathbb{R}^{2}$.
@@ 4. The volume of tubes
Let $M$ be a smooth (or real-analytic) sub-Riemannian manifold, equipped with a smooth (or real-analytic) measure $\mu$. Let $S\subset M$ be a bounded non-characteristic embedded submanifold of codimension $m\geq 1$, of class $C^{2}$ (without boundary) and with the extendibility property. Let $r_{0}=r_{0}(S)>0$ be its injectivity radius. Then the volume of the tube $r\mapsto\mu(T(r))$ is smooth (or real-analytic) on $0,r_{0})$. Furthermore, there exists a continuous density $\sigma_{m}$ on $A^{1}S$, defined by
@@
Let $\tilde{S}$ be the extension coming from the extendibility property, see 3.6. Thanks to the latter, we can cover $S$ with a finite number of charts where $S$ is the graph of a $C^{2}$ function. Thus, by a partition of unity argument, it is sufficient to consider the case in which $\mathrm{cl}(S)$ is contained in a single chart and is diffeomorphic to the closure of an open set of $\mathbb{R}^{n-m}$. More precisely, we assume there are smooth (or real-analytic) coordinates $(x,y)\in\mathbb{R}^{n-m}\times\mathbb{R}^{m}$ on an open neighbourhood $\mathcal{O}\subset M$ such that $\mathrm{cl}(S)\subset\mathcal{O}$, and there exists a $C^{2}$ function $h:\mathbb{R}^{n-m}\to\mathbb{R}^{m}$ and an open set $B\subset\mathbb{R}^{n-m}$ such that
@@
From the claim, it follows that $r\mapsto\mu(T(r))$ is smooth (or real-analytic) on $0,r_{0})$, and one can differentiate under the integral sign an infinite number of times in (87).
@@
so that the right hand side of (93) vanishes if kmk-m is odd, upon integration on $\mathbb{S}^{m-1}$.
@@
In the setting of Theorem 4.1, assume furthermore that $S$ is a two-sided hypersurface. Then the volume of the half-tube $r\mapsto\mu(T^{+}(r))$ is smooth (or real-analytic) on $0,r_{0})$. Moreover, $\mu(T^{+}(r))$ has the following Taylor expansion at $r=0$: